OpenStudy (anonymous):

If triangle RST is congruent to triangle NPQ which of the following is true angle r is congruent to angle P angle r is congruent to angle q angle t is congruent to angle q

OpenStudy (anonymous):

I chose angle r is congruent to angle q

OpenStudy (anonymous):

t is congruent to angle q

OpenStudy (anonymous):

Do you see why this is?

OpenStudy (anonymous):

not really I have issues when there isn't a picture

OpenStudy (anonymous):



OpenStudy (anonymous):

I cant draw but there you go

OpenStudy (anonymous):

haha thanks it does help :)

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

Given ΔRST is congruent to ΔNPQ, RT = 7x – 5, NQ = 5x + 11, find the length of Line of RTand Line of NQ. 8 51<-- 40 56

OpenStudy (anonymous):



OpenStudy (anonymous):

The answer is 8. You have to plug in the numbers and get the same answer.

OpenStudy (anonymous):

Thanks, Ugh I hate triangles! I can never seem to get them right.

OpenStudy (anonymous):

7(8)=56-5=51 5(8)=40+11=51

OpenStudy (anonymous):

Anything Else?

OpenStudy (anonymous):

ah see that's where I messed up, one more if you don't mind.

OpenStudy (anonymous):

this is the last one.

OpenStudy (anonymous):

I don't mind

OpenStudy (anonymous):

which of the following statements is true about the triangles below? ΔABC is congruent to ΔDEF by SSA ΔABC is congruent to ΔDEF by AAS<-- ΔABC is congruent to ΔDEF by SAS ΔABC is congruent to ΔDEF by ASA

OpenStudy (anonymous):

c

OpenStudy (anonymous):

because it shows a side, then an angle, then another side. Therefore SAS

OpenStudy (anonymous):

Thanks you were a real help!

OpenStudy (anonymous):

Anytime, if you ever need more help just fan me and ask for help